# Do graph neural network states contain graph properties?

### v2

@@ Abstract
Deep neural networks (DNNs) achieve state-of-the-art performance on many tasks, but this often requires increasingly larger model sizes, which in turn leads to more complex internal representations. Explainability techniques (XAI) have made remarkable progress in the interpretability of ML models. However, the non-relational nature of Graph neural networks (GNNs) make it difficult to reuse already existing XAI methods. While other works have focused on instance-based explanation methods for GNNs, very few have investigated model-based methods and, to our knowledge, none have tried to probe the embedding of the GNNs for well-known structural graph properties. In this paper we present a model agnostic explainability pipeline for GNNs employing diagnostic classifiers. This pipeline aims to probe and interpret the learned representations in GNNs across various architectures and datasets, refining our understanding and trust in these models.
@@ 1 Introduction
Graph Neural Networks (GNNs) are pivotal in harnessing non-Euclidean graph-structured data (Kipf & Welling, 2017) for tasks ranging from social network analysis to bioinformatics. Despite their success, the black-box nature of GNNs poses significant challenges as classical XAI methods can’t be directly applied on GNNs due to the lack of a regular structure (e.g. vertices can have different degrees). In this case, explaining a prediction means identifying important parts of the relational structure, or input features of nodes. An issue is that finding the explanation is itself a combinatorial problem, making XAI methods for GNN intractable (Longa et al., 2023a; Ying et al., 2019).
Previous surveys (Agarwal et al., 2023; Dai et al., 2022) highlighted the lack of comprehensive, robust and model-agnostic explainability methods. We also identified (sec C) few model-level explainability methods. They focus on explaining the decision-making process at a high level, often by generating graph patterns or motifs that influence the model’s predictions but none allow for the interpretability of intermediate layers and none highlight the role of graph properties. One paper identified the room for probing classifier (Akhondzadeh et al., 2023), as developed for Natural Language Processing (Giulianelli et al., 2018), (Belinkov, 2021). Their research question was ”will the hidden representation contain information about the number of hydrogen atoms or the presence of aromatic rings?”. We aim to address a more fundamental and general question : Does the hidden representation encode information about the graph-theoretic properties ? We propose a simple, model-agnostic, probing classifier pipeline to interpret GNN embeddings by probing for encoded graph properties (see fig. 1) across various architectures and datasets. We investigate both local properties like betweennes centrality, as well as global properties like average path length. To our knowledge, this is the first work to explore this direction.
We demonstrate the ability of diagnostic classifiers to effectively highlight graph-theoretic properties in GNN learned latent representations (fig. 7). We explore how different regularization techniques (none, $L_{2}$ weight decay, dropout) affect the representation of graph properties (fig. 14). We compare how various GNN architectures (GCN, R-GCN, GIN, GAT) differ in their ability to represent graph properties, analyzing whether these differences align with their mathematical frameworks (table 6). The approach is validated through applications to toxicity and fMRI datasets, confirming the alignment of probed properties with domain knowledge (table 8), before exploring the pipeline’s inferential power, uncovering structural properties that might not yet have been extensively studied (table 23).
Figure 1: Illustration of the probing pipeline. First, a GNN is trained on a specific task, such as classifying between smaller graphs (label 0) and larger graphs (label 1). Similar to how a CNN might organize images of similar shapes or textures into distinct regions in its feature space, the GNN embeddings might arrange graphs based on structural properties like their diameter. Next, we extract embeddings from the internal layers of the GNN. These embeddings are used to train the probing model—in this case, a binary classifier tasked with determining whether the embeddings encode predictive information about the graph’s diameter. If a linear probe has good performance ($R^{2}$ score) then there exists a hyperplane in the representation space that separates the inputs based on the property
@@ 2.1 Graph Neural Networks
Nowadays we have some theoretical understanding of the representational restrictions and capabilities of Graph Neural Networks (GNNs) with regard to the Weisfeiler-Lehman test (Akhondzadeh et al., 2023). We know that these cannot capture certain graph properties, such as connectivity or triangle-freeness (Franks et al., 2024; Kiefer, 2020; Kriege et al., 2018), due to their reliance on local structure. This constraint is also present in (message passing) GNNs.
Graph Convolutional Network (GCN) (Kipf & Welling, 2017) are GNNs where for a single layer, the node representation is computed as: $\bm{X}^{\prime}=\sigma\left(\tilde{\bm{D}}^{-1/2}\cdot\tilde{\bm{A}}\cdot\tilde{\bm{D}}^{-1/2}\cdot\bm{X}\cdot\bm{W}\right)$. We know that GNNs which rely solely on local information, like the GCN and its relational variant (R-GCN(Schlichtkrull et al., 2018), cannot compute important graph properties, such as girth and diameter or eigenvector centrality (Garg et al., 2020). We are therefore also investigating more globally aware networks like GAT (Graph Attention Network) (Veličković et al., 2018) and GIN (Graph Isomorphism Network) (Xu et al., 2019).
GAT makes use of self-attention and is thereby more expressive than the GCN. However, its reliance on feature-dependent weights and structure-free normalisation limits its ability to capture specific structural properties that do not directly depend on edges. This is particularly true for tasks where node features alone are not enough, and global graph structures are crucial (e.g., tasks requiring knowledge of subgraphs or non-local patterns). GIN aggregates node features in a way that mimics the Weisfeiler-Lehman test. By using the MLP equivalent to an Injective Update Function, GIN avoids oversimplifying the aggregation step, making it as expressive as the WL test. Thus, it is likely to excel at encoding complex graph properties and solving classification tasks.
@@ 2.2 Graph properties
Graph theory is a branch of mathematics that studies the properties and relationships of graphs. Graphs can be undirected or directed and analysed through both local and global properties. Local properties (like node degree or clustering coefficient) are based on the neighbors of a node. In contrast, global properties (such as diameter and characteristic path length) assess the overall graph structure. Global graph properties can be associated with higher level complex systems’ characteristics like the presence of repeated motifs in the graphs or information-flow properties. See the appendix B for a list of local and global properties used in our experiments.
We can distinguish different global properties, basic ones like the number of nodes a graph has, clustering and centrality ones, graph motifs and substructures, spectral and small-world properties. As an higher-order analysis, the recurrence of specific motifs within network substructures—such as triangles, cliques, or feed-forward loops can be seen as the fundamental building blocks that dictate the system’s functionality and resilience. Small-worldness, as characterised by Barabási (Albert & Barabási, 2002), reveal how networks can maintain short path lengths despite their expansive size and sparse connectivity. This kind of higher order properties are very interesting in order to understand how the macroscopic behaviour of complex systems emerges from the intricate interplay of their microscopic components (Barabási et al., 2002). For example how diseases spread in social networks, how neurons interact in the brain, or how information propagates through the Internet.
GNNs synthesise local topological features into global structures, abstract these representations into higher-order graph attributes. Each layer progressively expands the receptive field, mirroring how hierarchical feature learning works in convolutional neural networks (CNNs) for images. Probing their learnt representations should act as a scalable proxy to investigate how global arrangement and connectivity patterns influence a system’s function. In other terms, by dissecting these learned embeddings, we can possibly delve into the intricate relationships between a network’s macroscopic arrangement and its emergent behaviours. Based on the message passing paradigm in GNNs, as layers progress, one would expect an increased abstraction in the selection of graph properties. Initially, local features like node degree dominate, but deeper layers progressively capture more global properties, such as connectivity patterns and centrality. Through hierarchical pooling or readout mechanisms, GNNs can aggregate node embeddings into a single, global graph-level embedding. Graphs that share structural similarities or patterns of interaction among nodes are organised closely in the embedding space, allowing the model to differentiate between classes of graphs, such as those with and without long paths (fig. 1). This is why, AI engineers are likely to focus on the validation framework, emphasizing the alignment between probing results and the predictions derived from the message-passing paradigm—particularly insights revealed in the initial layers of the model. In contrast, from the standpoint of domain researchers—such as chemists or neuroscientists— the most compelling aspect is found in the later layers of the model, where the abstract representations become increasingly capable of rendering the problem linearly separable, thus facilitating clear interpretability of classification decisions and offering domain-specific interpretations.
@@ 2.3 Probing classifiers
In prior work (Hupkes et al., 2018) probing classifiers have been used for linguistic properties. Here, we adapt them for graph features. Unlike unsupervised techniques such as PCA or T-SNE, which are useful to visualize input data with regard to the embedding latent space, we adopt a supervised framework to quantitatively assess how specific properties are encoded within the embedding space of GNNs. Let $g:f_{l}(x)\mapsto\hat{z}$ represent a probing classifier, used to map the learned intermediate representations from the original model $f$ to a specific property $\hat{z}$. The choice of a linear classifier for $g$ is motivated primarily by its simplicity. If a linear probe performs well, it suggests the existence of a hyperplane in the representation space that separates the inputs based on their properties, indicating linear separability.
@@
Another advantage of a simple linear probe is avoiding the risk that a more complex classifier might infer features that are not actually used by the network itself (Hupkes et al., 2018). While other non-linear probes have been explored in the literature (Belinkov, 2021), even studies showing improved performance with complex probes maintain the same logic: $\operatorname{Perf}(g,f_{1},\mathcal{D}_{O},\mathcal{D}_{P})>\operatorname{Perf}(g,f_{2},\mathcal{D}_{O},\mathcal{D}_{P})$ holds across representations $f_{1}(x)$ and $f_{2}(x)$ when evaluated by a consistent probe $g$. This consistency ensures valid comparison, underscoring that if a property can be predicted well by a simple probe, it is likely relevant to the primary classification task. From an information-theoretic perspective, training the probing classifier $g$ can be viewed as estimating the mutual information between the learned representations $f_{l}(x)$ and the property $z$. This mutual information is denoted as $\mathrm{I}(\mathbf{z};\mathbf{h})$, where $\mathbf{z}$ refers to the property and $\mathbf{h}$ represents the intermediate representations (Belinkov, 2021).
@@
This supervised approach allows us to define hyperplanes or higher-dimensional decision boundaries that partition the embedding space according to the chosen graph property. The $R^{2}$ score serves as this information-theoretic measure indicating how well the hyperplane divides the inputs in the embedding space. The R2 score (, see appendix A for a formal definition) measures the proportion of variance in the dependent variable that is predictable from the independent variable(s). A $R^{2}$ near 1 indicates that the embeddings are highly informative about $\hat{z}$, suggesting that the neural model has internalized this property in a linearly accessible manner.
@@
By defining specific properties that could divide the embedding space and assessing how well the corresponding hyperplanes make the embedding space linearly separable, we gain quantitative insights into the abstract features aggregated within the embeddings. This method moves beyond mere hypothesis generation based on clustering patterns observed through techniques like PCA, providing a rigorous framework for understanding how well the embedding space represents complex graph properties. It can also be thought as complementary from the T-SNE and PCA visualisation techniques, as it provides a quantitative measure of the separability of the embeddings based on hypothesised properties of interest. We illustrate the evolution of the separability of graphs in the embeddings in figs. 4, 7 and 16 using a T-SNE visualisation and the corresponding separability with the properties thanks to the probing.
@@ 3 Datasets
All three datasets have the same setup: given a set of graphs $\left\{\mathcal{G}^{1},\mathcal{G}^{2},\ldots,\mathcal{G}^{N}\right\}$, predict the corresponding binary labels$\left\{y^{1},y^{2},\ldots,y^{N}\right\}$.
@@
inspired by (Agarwal et al., 2023) is designed to evaluate the compositionality of GNNs. It features two concepts: a 3x3 grid and a house-shaped graph made of five nodes. The dataset consists of Barabási-Albert (BA) graphs (Barabási, 2009) with a normal distribution of the number of nodes. The negative class includes a BA graph connected to either a grid or a house, while the positive class contains a BA graph connected to both a grid and a house (see fig. 2). In order to ensure that the average number of nodes is the same between classes, the number of nodes is a uniformly distributed between 6 and 21 for the grid graphs, between 7 and 22 for the house graphs, and between 1 and 16 when both are present. During generation, we ensure no test set leakage by removing isomorphisms. On 2,000 graphs, we perform an 80/20 train/test split.
For accurate classification, models need to identify and combine simple patterns. Recognizing isolated patterns or single node features is not sufficient. The dataset helps investigate how GNNs combine multiple concepts and addresses the “laziness” phenomenon, where networks learn patterns characterising only one class and predict the other by default (Longa et al., 2023b).
The dataset has been structured such that an optimal, linearly separable solution requires the combination of local properties, such as eigenvector centrality and betweenness centrality, or the identification of global structural motifs, like counting the number of squares (i.e., four-node cycles). A random Barabási-Albert graph can’t contain any four-node cycles, while a grid subgraph will consistently exhibit four such cycles. A house subgraph contains exactly one four-nodes cycle and one three-nodes cycle. Therefore, a graph that contains both a grid and a house will have a total of five four-node cycles. The presence of a three-node cycle could help the diagnostic of one type of graph in the negative class but is not necessary nor sufficient for solving the classification problem. On the contrary, counting the number of four-node cycle is necessary and sufficient. Thus, distinguishing between the classes does not really necessitates leveraging centrality-based measures but only recognizing the presence of a specific number of four-node cycles, enabling the model to effectively differentiate between the positive and negative classes. Thus the interesting results of fig. 4.
The Grid-House dataset serves a critical purpose in our study: it offers a controlled and well-defined environment for rigorously testing our hypotheses. The simplicity of these controlled constraints allow us to verify whether the GNN operates as intended in a setting where extraneous factors are minimized. If a model struggles with this dataset, it would likely underperform on real-world, more complex graphs, underscoring its diagnostic value.
@@
fMRI FC connectomes
consists of two parts. The Autism Brain Imaging Data Exchange I dataset contains 528 ASD patients and 571 typically developed (TD) individuals, the REST-meta-MDD dataset contains 848 MDD patients and 794 healthy controls. For both, the task is to classify these. We use the datasets with functional connectivity (FC) graphs, as prepared by  Zheng et al. (2023). We perform a 95/5% train-test random split.

### v3

@@ Abstract
Deep neural networks (DNNs) achieve state-of-the-art performance on many tasks, but this often requires increasingly larger model sizes, which in turn leads to more complex internal representations. Explainability techniques (XAI) have made remarkable progress in the interpretability of ML models. However, the non-euclidean nature of Graph Neural Networks (GNNs) makes it difficult to reuse already existing XAI methods. While other works have focused on instance-based explanation methods for GNNs, very few have investigated model-based methods and, to our knowledge, none have tried to probe the embedding of the GNNs for structural graph properties. In this paper we present a model agnostic explainability pipeline for Graph Neural Networks (GNNs) employing diagnostic classifiers. We propose to consider graph-theoretic properties as the features of choice for studying the emergence of representations in GNNs. This pipeline aims to probe and interpret the learned representations in GNNs across various architectures and datasets, refining our understanding and trust in these models.
@@ 1 Introduction
Graph Neural Networks (GNNs) are pivotal in harnessing graph-structured data (Kipf and Welling, 2017) for tasks ranging from social network analysis to bioinformatics. Despite their success, the black-box nature of GNNs poses significant challenges as classical XAI methods cannot be directly applied on GNNs due to the lack of a regular structure (e.g. vertices can have different degrees). In this case, explaining a prediction means identifying important parts of the relational structure, or input features of nodes. An issue is that finding the explanation is itself a combinatorial problem, making XAI methods for GNN intractable (Ying et al., 2019; Longa et al., 2023a).
The surveys by Dai et al. (2022) and Agarwal et al. (2023) highlighted the lack of comprehensive, robust and model-agnostic explainability methods. We also identified (see appendix C) few model-level explainability methods. Most works in GNN explainability present methods that identify a set of nodes and possibly their attributes as the explanation of a prediction (Vu and Thai, 2020; Zhang et al., 2021; Saha et al., 2022; Azzolin et al., 2023; Wang et al., 2023; Xuanyuan et al., 2023). Other works focus on explaining the decision-making process at a high level, often by generating graph patterns or motifs that influence the predictions (Ying et al., 2019; Yuan et al., 2020). However, oftentimes there is more information in the structure of the relationships between elements than in the juxtaposition of the elements themselves. Graph prediction ought to take into account structure that is invariant across the graph hierarchies and symmetries. Achieving this, we also get interpretability of intermediate layers, which previous methods do not provide.
One prior work identified the role probing classifiers could play (Akhondzadeh et al., 2023), as developed for Natural Language Processing (Giulianelli et al., 2018; Belinkov, 2021). That work focused on whether the hidden representation encoded the number of hydrogen atoms or the presence of aromatic rings. We aim to address the more fundamental and general question; finding out whether the hidden representation encodes information about the graph-theoretic properties. We propose graph theoretic properties as the best candidate for studying GNN inductive bias. Graph theoretic properties are natural algorithmic abstractions of graphs used in network science (Barabási et al., 2002). They would act as representational atoms (Bereska and Gavves, 2024), as features that GNNs would leverage to render graph classification problems linearly separable. The main hypothesis tested in this paper is whether GNNs leverage such algorithmic strategies to preserve, abstract, or discard structural information through the network’s layers. This work bridges GNN interpretability with the broader theory of linear feature representations in neural models (Nanda et al., 2023; Bereska and Gavves, 2024). We propose a model-agnostic pipeline to interpret GNN embeddings by probing for encoded graph properties (see fig. 1) across various architectures and datasets. We investigate both local properties like betweennes centrality, as well as global properties like average path length.  [^1] All code is publicly available from
https://github.com/TomPelletreauDuris/Probing-GNN-representations
Figure 1: Illustration of the probing pipeline. Note that graphs of label 0 have only grid or house while those of label 1 have both. If a linear probe has good performance ($R^{2}$ score) then there exists a hyperplane in the representation space that separates the inputs based on the property.
Our core contribution is that we show that using a diagnostic classifier, as illustrated in fig. 1, we can effectively highlight graph-theoretic properties in GNN learned latent representations (fig. 7). We further explore how different regularization techniques affect the representation of graph properties (see fig. 14). We also investigate how the GNN architecture (table 6) and datasets (table 8) affect the probing.
@@ 2.1 Graph Neural Networks
Graph Convolutional Network (GCN) (Kipf and Welling, 2017) are GNNs where for a single layer, the node representation is computed as: $\bm{X}^{\prime}=\sigma\left(\tilde{\bm{D}}^{-1/2}\cdot\tilde{\bm{A}}\cdot\tilde{\bm{D}}^{-1/2}\cdot\bm{X}\cdot\bm{W}\right)$. We know that GNNs which rely solely on local information, like the GCN and its relational variant (R-GCN(Schlichtkrull et al., 2018), cannot compute important graph properties, such as girth and diameter or eigenvector centrality (Garg et al., 2020). We are therefore also investigating more globally aware networks like GAT (Graph Attention Network) (Veličković et al., 2018) and GIN (Graph Isomorphism Network) (Xu et al., 2019). The models expressivity is based on the Weisfeiler-Lehman test (Akhondzadeh et al., 2023). GIN aggregates node features in a way that mimics the Weisfeiler-Lehman test. By using the MLP equivalent to an Injective Update Function, GIN avoids oversimplifying the aggregation step, making it as expressive as the WL test. Thus, it is likely to excel at encoding complex graph properties and solving classification tasks. GAT should theoretically be as expressive while we expect GAT to be slightly less expressive, GCN even less.
@@ 2.2 Graph properties
Graph theory is a branch of mathematics that studies the properties and relationships of graphs. Graphs can be undirected or directed and analysed through both local and global properties. Local properties (like node degree or clustering coefficient) are based on the neighbors of a node. In contrast, global properties (such as diameter and characteristic path length) assess the overall graph structure. Global graph properties can be associated with higher level complex systems’ characteristics like the presence of repeated motifs in the graphs or information-flow properties. See the appendix B for a list of local and global properties used in our experiments. We can distinguish different global properties, basic ones like the number of nodes a graph has, clustering and centrality ones, graph motifs and substructures, spectral and small-world properties. As an higher-order analysis, the recurrence of specific motifs within network substructures—such as triangles, cliques, or feed-forward loops can be seen as the fundamental building blocks that dictate the system’s functionality and resilience. Small-worldness, as characterised by Barabási (Albert and Barabási, 2002), reveal how networks can maintain short path lengths despite their expansive size and sparse connectivity. GNNs synthesise local topological features into global structures, abstract these representations into higher-order graph attributes. Each layer progressively expands the receptive field, aggregating local neighborhood properties that are relevant for the classification towards high level graph properties, mirroring how hierarchical feature learning works in convolutional neural networks (CNNs) for images. Through hierarchical pooling or readout mechanisms, GNNs can aggregate node embeddings into a single, global graph-level embedding. Based on the message passing paradigm in GNNs, as layers progress, one would expect an increased abstraction in the selection of graph properties. Initially, local features like node degree should dominate, but deeper layers progressively should capture more global properties, such as connectivity patterns and centrality. Graph-theoretic properties serve a symbolic role, offering interpretable, human-defined structure, while emergent features in GNNs reflect a connectionist process, arising through distributed representations shaped by task supervision.
@@ 2.3 Probing classifiers
In prior work (Hupkes et al., 2018) probing classifiers have been used for linguistic properties. The probing paradigm is a post-hoc explainability method Alain and Bengio (2018); Belinkov (2021). Probing is a concept-based approach in the larger field of mechanistic interpretability (Bereska and Gavves, 2024). Probing the learned representations for high-level concepts and patterns is a top-down approach to unravel a model’s decision-making process. It’s also called representation engineering (Zou et al., 2025). Here, we adapt them for graph features. Unlike unsupervised techniques such as PCA or T-SNE, which are useful to visualize input data with regard to the embedding latent space, we adopt a supervised framework to quantitatively assess how specific properties are encoded within the embedding space of GNNs. Let $g:f_{l}(x)\mapsto\hat{z}$ represent a probing classifier, used to map the learned intermediate representations from the original model $f$ to a specific property $\hat{z}$. The choice of a linear classifier for $g$ is motivated primarily by its simplicity. If a linear probe performs well, it suggests the existence of a hyperplane in the representation space that separates the inputs based on their properties, indicating linear separability.
The justification for simple linear probe comes from the linear representation hypothesis (Neal, 2017). The linear representation hypothesis proposes that features are directions in activation space, i.e. linear combinations of neurons (Bereska and Gavves, 2024). It’s well studied in LLMs (Park et al., 2025). Rather than assuming this hypothesis holds, linear probes can be taken as diagnostic tools to investigate whether learned representations organize information in linearly accessible ways for particular concepts and datasets.
@@
Another advantage of a simple linear probe is avoiding the risk that a more complex classifier might infer features that are not actually used by the network itself (Hupkes et al., 2018). While other non-linear probes have been explored in the literature (Belinkov, 2021), even studies showing improved performance with complex probes maintain the same logic: $\operatorname{Perf}(g,f_{1},\mathcal{D}_{O},\mathcal{D}_{P})>\operatorname{Perf}(g,f_{2},\mathcal{D}_{O},\mathcal{D}_{P})$ holds across representations $f_{1}(x)$ and $f_{2}(x)$ when evaluated by a consistent probe $g$. This consistency ensures valid comparison, underscoring that if a property can be predicted well by a simple probe, it is likely relevant to the primary classification task.
@@
Mathematically speaking, this supervised approach allows us to define hyperplanes or higher-dimensional decision boundaries that partition the embedding space according to the chosen graph property. The $R^{2}$ score serves as this information-theoretic measure indicating how well the hyperplane divides the inputs in the embedding space. The $R^{2}$ score (see appendix A for a formal definition) measures the proportion of variance in the dependent variable that is predictable from the independent variable(s). A $R^{2}$ near 1 indicates that the embeddings are highly informative about $\hat{z}$, suggesting that the neural model has internalized this property in a linearly accessible manner.
@@
By defining specific properties that could divide the embedding space and assessing how well the corresponding hyperplanes make the embedding space linearly separable, we gain quantitative insights into the abstract features aggregated within the embeddings. This method moves beyond mere hypothesis generation based on clustering patterns observed through techniques like PCA, providing a rigorous framework for understanding how well the embedding space represents complex graph properties. It can also be thought as complementary from the T-SNE and PCA visualisation techniques. It provides a quantitative measure of the separability of the embeddings based on the property of interest. Our best empirical observation of this is the correspondence between the T-SNE visualisation of embeddings and their corresponding $R^{2}$ scores in figs. 3, 7 and 4.
@@ 3 Datasets
[…]
@@
inspired by (Agarwal et al., 2023) is designed to evaluate the compositionality of GNNs. It features two concepts: a 3x3 grid and a house-shaped graph made of five nodes. The dataset consists of Barabási-Albert (BA) graphs (Barabási, 2009) with a normal distribution of the number of nodes. The negative class includes a BA graph connected to either a grid or a house, while the positive class contains a BA graph connected to both a grid and a house (see fig. 2). For accurate classification, models need to identify and combine simple patterns. Recognizing isolated patterns or single node features is not sufficient. In order to ensure that the classification can’t be solved by simply counting the number of nodes in a graph, the average number of nodes has to be the same between classes. The number of nodes is uniformly distributed between 6 and 21 for the grid graphs, between 7 and 22 for the house graphs, and between 1 and 16 when both are present. During generation, we ensure no test set leakage by removing isomorphisms. On 2,000 graphs, we apply an 80/20 train/test split. The dataset helps investigate how GNNs combine multiple concepts and addresses the “laziness” phenomenon, where networks learn patterns characterising only one class and predict the other by default (Longa et al., 2023b).
The dataset has been structured such that an optimal, linearly separable solution requires the identification of global structural motifs in the graph. A good strategy would be counting the number of squares (i.e., four-node cycles). Important note, a random Barabási-Albert graph cannot contain any four-node cycles. Meanwhile, a grid subgraph will consistently exhibit four such cycles, while a house subgraph contains exactly one four-node cycle and one three-node cycle. Therefore, a graph that contains both a grid and a house will have a total of five four-node cycles. The presence of a three-node cycle could help the diagnostic of one type of graph in the negative class but is not necessary nor sufficient for solving the classification problem. Conversely, counting the number of four-node cycles is necessary and sufficient. In our example in fig. 3 we see that a GIN architecture organizes the input graphs in its latent space layer by layer. At layer 5 (MLP1) it even differentiates grid-only from house-only graphs in the negative class (purple).
The Grid-House dataset serves a critical purpose in our study: it offers a controlled and well-defined environment for rigorously testing our hypotheses. By construction, we are sure the classification problem can be solved in an algorithmic way. The question becomes whether GNNs inductive bias converge to this algorithmic strategy. The simplicity of these controlled constraints allow us to verify whether the GNN operates as intended in a setting where extraneous factors are minimized.